# Authorship Obfuscation in Multilingual Machine-Generated Text Detection

### v2

@@ 1 Introduction
Recent advances in Language Modeling have birthed Large Language Models (LLMs) which exhibit significant improvements, including the ability to generate texts easily misconstrued as human-written (Zellers et al., 2019). In addition, LLMs have been found to exhibit emergent abilities (Wei et al., 2022), such as performing several NLP tasks at or over the human level, e.g., an LLM passing academic exams designed for humans in top 10% of test takers (Achiam et al., 2023). While this is a celebratory feat for the NLP community, LLMs currently have several shortcomings like the generation of hate and biased speech (Deshpande et al., 2023; Venkit et al., 2023), the generation of memorized and plagiarized content (Nasr et al., 2023; Carlini et al., 2021; Lee et al., 2023), and the generation of mis/disinformation (Lucas et al., 2023; Chen and Shu, 2023a; Vykopal et al., 2023; Zellers et al., 2019). Furthermore, we find that some of these LLMs like ChatGPT, Flan-T5 (Chung et al., 2022), Falcon (Almazrouei et al., 2023), LLaMA (Touvron et al., 2023), etc. can have their alignment tuning bypassed by jailbreaking prompting techniques (Lucas et al., 2023; Chen and Shu, 2023b). Therefore, due to the great potential to be misused, it is imperative that at the minimum, we can accurately distinguish Machine-Generated Texts (MGT) from human-written ones.
@@
(2) We provide a a unique dataset of multilingual obfuscated texts (human and machine-generated), which resulted into approximately 740k samples (although some of them contains the same text as original, i.e., the obfuscation failed). The dataset will be made publicly available upon the paper acceptance.
@@ 3 Methodology
The methodology is illustrated in Figure 1, summarizing key components. We benchmark 10 well-known AO methods attacking 37 MGT detection methods against texts in 11 languages (i.e., 10 $\times$ 37 $\times$ 11 = 4,070 combinations). All source codes, data, and full results will be publicly available to enable full replication of our study [^3] https://anonymous.4open.science/r/mAO.
@@ 3.2 Dataset
As a source of original unobfuscated texts, we use the MULTITuDE benchmark dataset (Macko et al., 2023b) containing about 8k human-written news texts along with 66k texts generated by 8 LLMs (i.e., alpaca-lora, gpt-3.5-turbo, gpt-4, llama, opt, opt-iml-max, text-davinci-003, vicuna) in 11 languages (i.e., Arabic, Catalan, Chinese, Czech, Dutch, English, German, Portuguese, Russian, Spanish, Ukrainian). For each AO method, we duplicated the MULTITuDE data and replaced the texts by their obfuscated version (we retry this if some obfuscated texts are the same as the unobfuscated ones; we limit the number of trials to 10). We evaluated similarity between the obfuscated and the original texts using a range of similarity metrics along with a human check of data pseudo-random subset to gauge the quality of the obfuscation and detect potential pitfalls (see Appendix C).
@@ 3.3 MGT Detection Methods
As pre-trained methods, we include publicly available detection models that were fine-tuned for MGT detection tasks using different datasets (i.e., out-of-distribution data) and use them in a zero-shot manner in this work. These include RoBERTa-Base and RoBERTa-Large OpenAI Detectors (Solaiman et al., 2019), ChatGPT-detector-RoBERTa and ChatGPT-detector-RoBERTa-Chinese (Guo et al., 2023a), Longformer Detector (Li et al., 2023), RoBERTa-base-autextification-Detection555www.huggingface.co/arincon/roberta-base-autextification, and ruRoBERTa-ruatd-binary [^6] https://huggingface.co/orzhan/ruroberta-ruatd-binary.
@@
As for statistical methods, we include single-metric methods (Entropy by Lavergne et al., 2008, LogLikelihood Mitchell et al., 2023, Rank Gehrmann et al., 2019, LogRank Mitchell et al., 2023, LLMDeviation by Wu and Xiang, 2023, DetectLLM-LRR by Su et al., 2023), multi-metric methods (GLTR Test 2 Features by Gehrmann et al., 2019, MFD by Wu and Xiang, 2023), and perturbation-based methods (DetectGPT by Mitchell et al., 2023, DetectLLM-NPR by Su et al., 2023). Due to the long run-time of perturbation-based methods, these have been used only on a portion of the test set. All of the mentioned statistical methods are implemented in the publicly available IMGTB framework [^7] https://github.com/kinit-sk/IMGTB (Spiegel and Macko, 2023), which was used in its default configuration (i.e., Logistic Regression classifier for predictions without hyper-parameters optimization). Analogously to the fine-tuned methods, the classifiers of the statistical methods have been trained using the original MULTITuDE data only (i.e., unobfuscated).
Rank MGT Detection Method Category AUC ROC Macro avg. Macro avg. Macro avg.
F1-score F1-score (optimal) F1-score (5% FPR)
1 XLM-RoBERTa-large (all) F 0.9247 0.5745 0.5119 0.4538
18 MFD S 0.6713 0.4799 0.3564 0.2526
19 RoBERTa-large-OpenAI-Detector P 0.6618 0.2266 0.4972 0.4997
Table 2: Detection performance comparison of MGT detection methods using all MULTITuDE test data (human and machine samples) and machine samples obfuscated by each AO method showing the best-ranked method for each category (F refers to fine-tuned, P to pre-trained, and S to statistical). The full version available in Table 13.
@@ 4 Experiments and Results
To compare MGT detection performance, we use standard metrics, namely AUC ROC (area under the curve of receiver operating characteristic) as a classification-threshold independent metric (not affected by a threshold calibration on a domain data) and Macro avg. F1-score as a metric balancing between a precision and a recall of the classification commonly used for the MGT detection task using imbalanced datasets. We further provide Macro avg. F1-score values theoretically achievable on the used data with the classification thresholds calibrated based on the ROC curve for optimal (i.e., maximal difference between TPR and FPR, where TPR is true positive rate and FPR is false positive rate) and 5% FPR conditions. We are aware that such a performance is not actually achievable due to calibration on the evaluation data (i.e., data leakage). Therefore, we are not using it to compare performance of the MGT detection methods, but to evaluate effect of AO on the detection (i.e., compare AO methods), while not being biased due to sub-optimal predictions.
@@
Before diving deeper into the experiments, let us briefly examine the general detection performance on our data (i.e., difficulty of the data for this task). Table 2 shows the best performing MGT detection methods for each category evaluated using the whole test set (i.e., the original MULTITuDE human and machine texts from the test split and the obfuscated machine texts); for the comparison of all 37 detectors, see Table 13 in Appendix E. Obfuscated human texts are not used in the experiments due to uncertain labeling after obfuscation (i.e., whether the texts modified by machine should still be considered as human) (Tripto et al., 2024). Based on the results of achieved AUC ROC values, existing detection methods can be also used for distinguishing between human-written and obfuscated machine-generated texts. However, based on Macro avg. F1-score maximum values, the data are quite challenging for the task. A high difference between AUC ROC and Macro avg. F1-score values are due to highly imbalanced data with a ratio between human and machine classes around 1:80.
@@ 4.1 Multilingual Capability of AO Methods
Backtranslation and other paraphrasers are more successful than ChatGPT. However, we have used only a basic prompt for paraphrasing by ChatGPT. A more sophisticated prompt directing the writing style could help to increase the ASR. Nevertheless, the ChatGPT paraphrasing does not suffer by language change in multilingual settings as do the English-only dedicated paraphrasers (Pegasus-parphrase and DIPPER).
Human validation of readability. Such an experiment properly evaluates the impact of AO methods on the automated MGT detection methods. However, the AO methods could have severely damaged the texts, which would be either unreadable by humans or clearly suspicious. Therefore, we have done a manual human check (see Appendix C) of the obfuscated data sample by three human evaluators (average Pearson correlation between annotations of 0.7, total annotation agreement accuracy of 0.65 with majority 2 of 3 accuracy of 0.99). Based on such a quality check along with text-similarity analysis between unobfuscated and obfuscated texts, Pegasus-paraphrase and DIPPER AO methods are unusable in the multilingual settings. Therefore, these two methods are disqualified from the further experiments (although included in the Appendices for a reference).
@@ Acknowledgements
Part of the research results was obtained using the computational resources procured in the national project National competence centre for high performance computing (project code: 311070AKF2) funded by European Regional Development Fund, EU Structural Funds Informatization of Society, Operational Program Integrated Infrastructure.
@@ C.1 Human Quality Check
The aim of the human quality check is to evaluate the obfuscated data quality, i.e., whether the obfuscated data is/is not clearly damaged from a human view. Massive human study is infeasible due to the scope and complexity of the work (it would require proficient annotators in 11 languages); therefore, a manual quality check by three human evaluators (selected from the paper authors) has been executed with a purpose of identifying unusable AO methods.
@@ Appendix E Results Data
F1-score F1-score (optimal) F1-score (1% FPR) F1-score (5% FPR)
1 XLM-RoBERTa-large (all) F 0.9247 0.5745 0.5119 0.2296 0.4538
2 XLM-RoBERTa-large (ru) F 0.9231 0.5983 0.4709 0.3298 0.4290
3 mDeBERTa-v3-base (all) F 0.9076 0.5388 0.4917 0.2826 0.4057
4 mDeBERTa-v3-base (ru) F 0.8895 0.6434 0.5160 0.1497 0.3405
5 mDeBERTa-v3-base (es) F 0.8616 0.5185 0.4743 0.1633 0.3573
7 mGPT (all) F 0.8511 0.5347 0.4640 0.2525 0.3110
9 mGPT (ru) F 0.8427 0.5901 0.4924 0.0110 0.0110
10 XLM-RoBERTa-large (es) F 0.8346 0.5035 0.4996 0.0110 0.2747
11 mGPT (es) F 0.8312 0.5074 0.4363 0.0110 0.3217
12 OPT-IML-Max-1.3B (all) F 0.8261 0.5265 0.4406 0.0110 0.1958
13 OPT-IML-Max-1.3B (es) F 0.7697 0.5024 0.4905 0.0110 0.0110
16 OPT-IML-Max-1.3B (ru) F 0.7101 0.5346 0.4063 0.0776 0.1961
17 XLM-RoBERTa-large (en) F 0.6815 0.5285 0.3734 0.1274 0.2073
18 MFD S 0.6713 0.4799 0.3564 0.1069 0.2526
20 Entropy S 0.6191 0.4972 0.2433 0.1562 0.2229
21 mGPT (en) F 0.6178 0.4936 0.3181 0.0110 0.0110
22 Longformer Detector P 0.6135 0.4972 0.2531 0.0582 0.1362
23 mDeBERTa-v3-base (en) F 0.6112 0.4660 0.4166 0.0152 0.0775
25 OPT-IML-Max-1.3B (en) F 0.5824 0.5452 0.1489 0.0968 0.1489
26 DetectLLM-NPR S 0.5764 0.4926 0.2844 0.0636 0.1469
28 GLTR Test 2 S 0.5385 0.4922 0.3001 0.0454 0.1188
29 DetectGPT S 0.5382 0.4926 0.2700 0.0581 0.1231
30 ChatGPT-Detector-RoBERTa P 0.5311 0.1036 0.4972 0.0110 0.4551
31 DetectLLM-LRR S 0.5250 0.4966 0.2587 0.0573 0.1494
33 LogRank S 0.4669 0.4965 0.2504 0.0344 0.0635
34 LLMDeviation S 0.4589 0.4967 0.2349 0.0292 0.0664
35 LogLikelihood S 0.4508 0.4966 0.2521 0.0323 0.0595
36 ruRoBERTa-ruatd-binary P 0.4406 0.4772 0.0110 0.0110 0.0110
37 Rank S 0.3859 0.4972 0.0110 0.0110 0.0110
Table 13: Detection performance comparison of all MGT detection methods using all MULTITuDE test data (human and machine samples) and machine samples obfuscated by each AO method. Macro avg. F1-score (optimal) is just a theoretical performance and refers to this metric after calibration of the classification threshold based on ROC curve maximizing difference between TPR (true positive rate) and FPR (false positive rate). Similarly, Macro avg. F1-score (1% FPR) and (5% FPR) refer to this metric for the classification threshold closest to (but below) 1% and 5% of FPR. F refers to fine-tuned, P to pre-trained, and S to statistical MGT detection methods category.
@@
Table 13 shows a comparison of detection performance of all MGT detection methods using all MULTITuDE test data (human and machine samples) and machine samples obfuscated by each AO method. Table 14 shows a per-language comparison of AUC ROC detection performance of all MGT detection methods using subset of MULTITuDE test data, namely all human samples along with the machine samples generated by the selected three best LLM generators (gpt-4, gpt-3.5-turbo, vicuna), generating the texts most similar to human counterparts based on automated similarity metrics. Tables 1517 include resulted mean values (across detection methods) of AUC ROC detection performance per each AO method data and per each test language. To see differences between MGT detection method categories, we provide the results separately for them. Furthermore, we provide AUC ROC drop in Tables 1820, reflecting the robustness of MGT detection method categories in regard to each AO method.
@@
Absolute AUC ROC $\uparrow$ Relative AUC ROC Difference $\downarrow$
AO Method (Category)
originally trained
all
m2m100-1.2B
ChatGPT
Homoglyph Attack
DFTFooler
all
m2m100-1.2B
ChatGPT
Homoglyph Attack
DFTFooler

### v3

@@ 1 Introduction
Recent advances in Language Modeling have birthed Large Language Models (LLMs) which exhibit significant improvements, including the ability to generate texts easily misconstrued as human-written (Zellers et al., 2019). In addition, LLMs have been found to exhibit emergent abilities (Wei et al., 2022), such as performing several NLP tasks at or over the human level, e.g., an LLM passing academic exams designed for humans in top 10% of test takers (Achiam et al., 2023). While this is a celebratory feat for the NLP community, LLMs currently have several shortcomings like the generation of hate and biased speech (Deshpande et al., 2023; Venkit et al., 2023), the generation of memorized and plagiarized content (Nasr et al., 2023; Carlini et al., 2021; Lee et al., 2023), and the generation of mis/disinformation (Lucas et al., 2023; Chen and Shu, 2023a; Vykopal et al., 2024; Zellers et al., 2019). Furthermore, we find that some of these LLMs like ChatGPT, Flan-T5 (Chung et al., 2022), Falcon (Almazrouei et al., 2023), LLaMA (Touvron et al., 2023), etc. can have their alignment tuning bypassed by jailbreaking prompting techniques (Lucas et al., 2023; Chen and Shu, 2023b). Therefore, due to the great potential to be misused, it is imperative that at the minimum, we can accurately distinguish Machine-Generated Texts (MGT) from human-written ones.
@@
(2) We provide a unique dataset of multilingual obfuscated texts (human and machine-generated), which resulted into approximately 740k samples (although some of them contains the same text as original, i.e., the obfuscation failed). The dataset is available at Zenodo upon request for research purposes only [^1] https://doi.org/10.5281/zenodo.13846588.
@@ 3 Methodology
The methodology is illustrated in Figure 1, summarizing key components. We benchmark 10 well-known AO methods attacking 37 MGT detection methods against texts in 11 languages (i.e., 10 $\times$ 37 $\times$ 11 = 4,070 combinations). All source codes, data, and full results are publicly available to enable full replication of our study [^4] https://github.com/kinit-sk/mAO.
@@ 3.2 Dataset
As a source of original unobfuscated texts, we use the MULTITuDE benchmark dataset (Macko et al., 2023b) containing about 8k human-written news texts along with 66k texts generated by 8 LLMs (i.e., alpaca-lora, gpt-3.5-turbo, gpt-4, llama, opt, opt-iml-max, text-davinci-003, vicuna) in 11 languages (i.e., Arabic, Catalan, Chinese, Czech, Dutch, English, German, Portuguese, Russian, Spanish, Ukrainian). For each AO method, we duplicate the MULTITuDE data and replace the texts by their obfuscated version (we retry this if some obfuscated texts are the same as the unobfuscated ones; we limit the number of trials to 10). We evaluate similarity between the obfuscated and the original texts using a range of similarity metrics along with a human check of data pseudo-random subset to gauge the quality of the obfuscation and detect potential pitfalls (see Appendix C).
@@ 3.3 MGT Detection Methods
As pre-trained methods, we include publicly available detection models that were fine-tuned for MGT detection tasks using different datasets (i.e., out-of-distribution data) and use them in a zero-shot manner in this work. These include RoBERTa-Base and RoBERTa-Large OpenAI Detectors (Solaiman et al., 2019), ChatGPT-detector-RoBERTa and ChatGPT-detector-RoBERTa-Chinese (Guo et al., 2023a), Longformer Detector (Li et al., 2023), RoBERTa-base-autextification-Detection666www.huggingface.co/arincon/roberta-base-autextification, and ruRoBERTa-ruatd-binary [^7] https://huggingface.co/orzhan/ruroberta-ruatd-binary.
@@
As for statistical methods, we include single-metric methods (Entropy by Lavergne et al., 2008, LogLikelihood by Mitchell et al., 2023, Rank by Gehrmann et al., 2019, LogRank by Mitchell et al., 2023, LLMDeviation by Wu and Xiang, 2023, DetectLLM-LRR by Su et al., 2023), multi-metric methods (GLTR Test 2 Features by Gehrmann et al., 2019, MFD by Wu and Xiang, 2023), and perturbation-based methods (DetectGPT by Mitchell et al., 2023, DetectLLM-NPR by Su et al., 2023). Due to the long run-time of perturbation-based methods, these have been used only on a portion of the test set. All of the mentioned statistical methods are implemented in the publicly available IMGTB framework [^8] https://github.com/kinit-sk/IMGTB (Spiegel and Macko, 2024), which was used in its default configuration (i.e., Logistic Regression classifier for predictions without hyper-parameters optimization). Analogously to the fine-tuned methods, the classifiers of the statistical methods have been trained using the original MULTITuDE data only (i.e., unobfuscated).
(sorted) F1-score F1-score (optimal) F1-score (1% FPR) F1-score (5% FPR)
1 XLM-RoBERTa-large (all) F 0.9247 0.5745 0.5119 0.2296 0.4538
2 XLM-RoBERTa-large (ru) F 0.9231 0.5983 0.4709 0.3298 0.4290
3 mDeBERTa-v3-base (all) F 0.9076 0.5388 0.4917 0.2826 0.4057
4 mDeBERTa-v3-base (ru) F 0.8895 0.6434 0.5160 0.1497 0.3405
5 mDeBERTa-v3-base (es) F 0.8616 0.5185 0.4743 0.1633 0.3573
7 mGPT (all) F 0.8511 0.5347 0.4640 0.2525 0.3110
9 mGPT (ru) F 0.8427 0.5901 0.4924 0.0110 0.0110
10 XLM-RoBERTa-large (es) F 0.8346 0.5035 0.4996 0.0110 0.2747
11 mGPT (es) F 0.8312 0.5074 0.4363 0.0110 0.3217
12 OPT-IML-Max-1.3B (all) F 0.8261 0.5265 0.4406 0.0110 0.1958
13 OPT-IML-Max-1.3B (es) F 0.7697 0.5024 0.4905 0.0110 0.0110
16 OPT-IML-Max-1.3B (ru) F 0.7101 0.5346 0.4063 0.0776 0.1961
17 XLM-RoBERTa-large (en) F 0.6815 0.5285 0.3734 0.1274 0.2073
18 MFD S 0.6713 0.4799 0.3564 0.1069 0.2526
20 Entropy S 0.6191 0.4972 0.2433 0.1562 0.2229
21 mGPT (en) F 0.6178 0.4936 0.3181 0.0110 0.0110
22 Longformer Detector P 0.6135 0.4972 0.2531 0.0582 0.1362
23 mDeBERTa-v3-base (en) F 0.6112 0.4660 0.4166 0.0152 0.0775
25 OPT-IML-Max-1.3B (en) F 0.5824 0.5452 0.1489 0.0968 0.1489
26 DetectLLM-NPR S 0.5764 0.4926 0.2844 0.0636 0.1469
28 GLTR Test 2 S 0.5385 0.4922 0.3001 0.0454 0.1188
29 DetectGPT S 0.5382 0.4926 0.2700 0.0581 0.1231
30 ChatGPT-Detector-RoBERTa P 0.5311 0.1036 0.4972 0.0110 0.4551
31 DetectLLM-LRR S 0.5250 0.4966 0.2587 0.0573 0.1494
33 LogRank S 0.4669 0.4965 0.2504 0.0344 0.0635
34 LLMDeviation S 0.4589 0.4967 0.2349 0.0292 0.0664
35 LogLikelihood S 0.4508 0.4966 0.2521 0.0323 0.0595
36 ruRoBERTa-ruatd-binary P 0.4406 0.4772 0.0110 0.0110 0.0110
37 Rank S 0.3859 0.4972 0.0110 0.0110 0.0110
Table 2: Detection performance comparison of all MGT detection methods using all MULTITuDE test data (human and machine samples) and machine samples obfuscated by each AO method. Macro avg. F1-score (optimal) is just a theoretical performance and refers to this metric after calibration of the classification threshold based on ROC curve maximizing difference between TPR (true positive rate) and FPR (false positive rate). Similarly, Macro avg. F1-score (1% FPR) and (5% FPR) refer to this metric for the classification threshold closest to (but below) 1% and 5% of FPR. For the exact calculations of such thresholds, see the published source code. F refers to fine-tuned, P to pre-trained, and S to statistical MGT detection methods category.
@@ 4 Experiments and Results
To compare MGT detection performance, we use standard metrics, namely AUC ROC (area under the curve of receiver operating characteristic) as a classification-threshold independent metric (not affected by a threshold calibration on a domain data) and Macro avg. F1-score as a metric balancing between a precision and a recall of the classification commonly used for the MGT detection task using imbalanced datasets. We further provide Macro avg. F1-score values theoretically achievable on the used data with the classification thresholds calibrated based on the ROC curve for optimal (i.e., maximal difference between TPR and FPR, where TPR is true positive rate and FPR is false positive rate), 1% FPR and 5% FPR conditions. We are aware that such a performance is not actually achievable due to calibration on the evaluation data (i.e., data leakage). Therefore, we are not using it to compare performance of the MGT detection methods, but to evaluate effect of AO on the detection (i.e., compare AO methods), while not being biased due to sub-optimal predictions.
@@
Before diving deeper into the experiments, let us briefly examine the general detection performance on our data (i.e., difficulty of the data for this task). Table 2 shows a comparison of detection performance of all MGT detection methods using the whole test set (i.e., the original MULTITuDE human and machine texts from the test split and the obfuscated machine texts). Obfuscated human texts are not used in the experiments due to uncertain labeling after obfuscation (i.e., whether the texts modified by machine should still be considered as human) (Tripto et al., 2024). Based on the results of achieved AUC ROC values, existing detection methods can be also used for distinguishing between human-written and obfuscated machine-generated texts. However, based on Macro avg. F1-score maximum values, the data are quite challenging for the task. A high difference between AUC ROC and Macro avg. F1-score values are due to highly imbalanced data with a ratio between human and machine classes around 1:80.
@@ 4.1 Multilingual Capability of AO Methods
Backtranslation and other paraphrasers are more successful than ChatGPT in obfusation. However, we have used only a basic prompt for paraphrasing by ChatGPT. A more sophisticated prompt directing the writing style could help to increase the ASR. Nevertheless, the ChatGPT paraphrasing does not suffer by language change in multilingual settings as do the English-only dedicated paraphrasers (Pegasus-parphrase and DIPPER).
@@
4.1.1 Human Validation of Readability
Such an experiment properly evaluates the impact of AO methods on the automated MGT detection methods. However, the AO methods could have severely damaged the texts, which would be either unreadable by humans or clearly suspicious. Therefore, we have done a manual human check (see Appendix C) of the obfuscated data sample by three human evaluators (average Pearson correlation between annotations of 0.7, total annotation agreement accuracy of 0.65 with majority 2 of 3 accuracy of 0.99). Based on such a quality check along with text-similarity analysis between unobfuscated and obfuscated texts, Pegasus-paraphrase and DIPPER AO methods are unusable in the multilingual settings. Therefore, these two methods are disqualified from the further experiments (although included in the Appendices for a reference).
@@ 4.3 Multilingual Obfuscation for Data Augmentation
Absolute AUC ROC $\uparrow$ Relative AUC ROC Difference $\downarrow$
AO Method (Category)
originally trained
all
m2m100-1.2B
ChatGPT
Homoglyph Attack
DFTFooler
all
m2m100-1.2B
ChatGPT
Homoglyph Attack
DFTFooler
@@ Acknowledgements
Part of the research results was obtained using the computational resources provided by the CloudBank (https://www.cloudbank.org/), which was supported by the NSF award #1925001, and those procured in the national project National competence centre for high performance computing (project code: 311070AKF2) funded by European Regional Development Fund, EU Structural Funds Informatization of Society, Operational Program Integrated Infrastructure.
@@ C.1 Human Quality Check
The aim of the human quality check is to evaluate the obfuscated data quality, i.e., whether the obfuscated data is/is not clearly damaged from a human view. Massive human study is infeasible due to the scope and complexity of the work (it would require proficient annotators in 11 languages); therefore, a manual quality check by three human evaluators (selected from the paper authors) has been executed to identify unusable AO methods.
@@ Appendix E Results Data
Table 14 shows a per-language comparison of AUC ROC detection performance of all MGT detection methods using subset of MULTITuDE test data, namely all human samples along with the machine samples generated by the selected three best LLM generators (gpt-4, gpt-3.5-turbo, vicuna), generating the texts most similar to human counterparts based on automated similarity metrics. Tables 1517 include resulted mean values (across detection methods) of AUC ROC detection performance per each AO method data and per each test language. To see differences between MGT detection method categories, we provide the results separately for them. Furthermore, we provide AUC ROC drop in Tables 1820, reflecting the robustness of MGT detection method categories in regard to each AO method.